An ambulance blows its 160-Hz horn while traveling at 40.0 m/s in still air on a day when the speed of sound is 340 m/s. What frequency is observed by the train’s driver on the train?


1
Expert's answer
2021-06-21T11:41:21-0400

We can find the frequency observed by the train’s driver on the train from the Doppler Shift formula. Unfortunately, from the definition of the question we don't know the direction of the motion of the train and its speed. So, let's suppose that the train moves towards the ambulance at 60 m/s. Then, we get:


"f_o=f_s(\dfrac{v+v_o}{v-v_s}),""f_o=160\ Hz\cdot(\dfrac{340\ \dfrac{m}{s}+60\ \dfrac{m}{s}}{340\ \dfrac{m}{s}-40\ \dfrac{m}{s}})=213\ Hz."